BrotherBear:

I think i may have already asked for help with this but i couldnt find it. please help- (screenshots included somewhere)

BrotherBear:

BrotherBear:

MiraAngel:

@jhonyy9

MiraAngel:

and... https://questioncove.com/users/brotherbear#/updates/600735c24a24f0eb0abc5c00 There's the other one. An easy way to find old questions you posted is to go to your account and click questions asked.

BrotherBear:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @MiraAngel and... https://questioncove.com/users/brotherbear#/updates/600735c24a24f0eb0abc5c00 There's the other one. An easy way to find old questions you posted is to go to your account and click questions asked. \(\color{#0cbb34}{\text{End of Quote}}\) ohhh ok ty.

MiraAngel:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BrotherBear \(\color{#0cbb34}{\text{Originally Posted by}}\) @MiraAngel and... https://questioncove.com/users/brotherbear#/updates/600735c24a24f0eb0abc5c00 There's the other one. An easy way to find old questions you posted is to go to your account and click questions asked. \(\color{#0cbb34}{\text{End of Quote}}\) ohhh ok ty. \(\color{#0cbb34}{\text{End of Quote}}\) Your Welcome! Good luck! :)

jhonyy9:

so just use the Pythagora's theorem c^2 = a^2 +b^2 c hypothenuse a,b legs

BrotherBear:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 so just use the Pythagora's theorem c^2 = a^2 +b^2 c hypothenuse a,b legs \(\color{#0cbb34}{\text{End of Quote}}\) so would the answer be D?

jhonyy9:

16^2 = 8^2 +x^2 x^2 = 16^2 -8^2 x = sqrt 192 = 8sqrt3 yes D.